Hi- my class is working on Equestions with Variables/ Decimals/ Fractions. I honestly do not understand any of it. The problem I am working on is : 2a + 4.2 = 3a - 1.6 . I have tried to icolate the variable , but I still am stuck. I also do understand the decimals.
Please help me.
Sincerely, Sebastian Mosier

Answer by checkley71(8403)

2A+4.2=3A-1.6 FIRST YOU NEED TO COMBINE THE UNKNOWN TERMS(A) AND ALSO COMBINE THE NUMBERS.
TO DO THIS YOU SUBTRACT (-2A) FROM BOTH SIDES OF THE EQUATION WHICH GIVES
2A-2A+4.2=3A-2A-1.6 NOW WE HAVE ELIMINATED THE A FROM THE LEFT SIDE OF THE EQUATION AND HAVE 4.2=A-1.6 NOW WE ADD 1.6 TO BOTH SIDES AND WE GET
4.2+1.6=A OR 5.8=A
PROOF SUBSTITUTE 5.8 FOR A IN THE ORIGINAL EQUATION WE HAVE
2(5.8)+4.2=3(5.8)-1.6 OR 11.6 +4.2=17.4-1.6 OR 15.8=15.8
REMEMBER YOU CAN DO ANYTHING (ADD, SUBTRACT, MULTIPLY, DIVIDE, RAISE TO A POWER OR TAKE THE ROOT OF ANY VALUE) AS LONG AS YOU PERFORM THIS OPERATIOPN TO BOTH SIDES OF THE EQUATION. IT'S A LOGICAL BALANCE BEAM THEREFORE YOU MUST CHANGE BOTH SIDES OF THE EQUATION IN ORDER TO NOT CHANGE THE BALANCE.